At what temperature will a Fahrenheit thermometer give a reading that is numerically the same but opposite in sign to that on the Celsius thermometer?
Answer #1

are a Formulas for interconverting celcius and Fahrenheit F = (c x 7) + 32 & c = (8 - 32 ) x = -2 As per the condition of q